# DecompGAIL: Learning Realistic Traffic Behaviors with Decomposed Multi-Agent Generative Adversarial Imitation Learning

### v1

@@ Results.
Table 1: Results on the WOSAC 2025 leaderboard test split. Waymo (2025)
SMART-tiny-DecompGAIL (ours) 0.7864 0.4919 0.8152 0.9176 1.4209
SMART-R1 Pei et al. (2025) 0.7858 0.4944 0.8110 0.9201 1.2885
SMART-tiny-RLFTSim Ahmadi & Schofield (2025) 0.7857 0.4927 0.8129 0.9183 1.3252
TrajTok Zhang et al. (2025) 0.7852 0.4887 0.8116 0.9207 1.3179
SMART-tiny-CLSFT Zhang et al. (2024) 0.7846 0.4931 0.8106 0.9177 1.3065
UniMM Lin et al. (2025) 0.7829 0.4914 0.8089 0.9161 1.2949
SMART-tiny Wu et al. (2024) 0.7814 0.4854 0.8089 0.9153 1.3931
InfGen Peng et al. (2025) 0.7731 0.4493 0.8084 0.9127 1.4252
@@ 5.5 Qualitative results
We provide qualitative demonstrations of our fine-tuned model in fig.˜4, showcasing one rollout in two distinct scenarios. In both cases, the policy generates realistic multi-agent trajectories, maintaining collision-free and on-road behavior throughout the full 8-second horizon. This highlights the ability of our model to sustain expert-like motion even in complex interactive settings.
To further interpret model behavior, we visualize the per-agent rewards estimated by our decomposed discriminator, encoded as the color of each agent at every timestep. As illustrated, the discriminator consistently assigns lower rewards to agents that deviate from realistic behavior—for example, when a vehicle approaches a near-collision situation or begins drifting toward the road edge. Conversely, agents following smooth and expert-like motion patterns are assigned high rewards. Additional rollout examples, including more challenging interactive scenarios, are provided in the supplementary videos.
@@ 6 Conclusion
7 Reproducibility statement
We have taken multiple steps to ensure the reproducibility of our work. Our dataset and pretrained backbone are the same as those provided in the open-source CAT-K repository (https://github.com/NVlabs/catk/tree/main). Detailed descriptions of the model architecture and all hyperparameters are included in Section A and in the appendix. We also report the official performance of our method on the public WOSAC leaderboard, which enables transparent comparison against other approaches. We plan to release our code in the future to further facilitate reproducibility.
@@ Appendix A Implementation Details
We pretrain the policy network with BC for 32 epochs following CAT-K [46], taking about 44 hours, and then fine-tune with DecompGAIL for 2 additional epochs, taking about 10 hours. During fine-tuning, the map encoder is frozen to reduce GPU memory usage and the policy network and discriminator network share the map encoder, enabling larger batch sizes with negligible performance loss. Both pretraining and fine-tuning are conducted on 8 $\times$ H800 80GB GPUs with a total batch size of 80. We use Adam with weight decay of 0.01 for optimization, with a policy learning rate of $5\times 10^{-4}$ during pretraining and $5\times 10^{-5}$ during fine-tuning, and a discriminator learning rate of $1\times 10^{-4}$. To stabilize adversarial training, we generate fresh rollouts at each training step using the current policy, then update the discriminator and policy once. Additional hyperparameters for DecompGAIL fine-tuning are provided in table˜3.
@@
$\beta$ in $\lambda_{ij}$ 10
Appendix B Ethics statement
@@
Appendix C The Use of Large Language Models (LLMs)

### v2

@@ 5.4 Ablation (Q3)
5.5 Hyperparameter
In this section, we analyze the effect of the decay parameters $\alpha$ and $\beta$ used for interaction weighting and social rewards on the realism meta-metric. We first conduct a sweep over decay scales $\alpha\in\{2.5,5,10,20\}$ and decay ranges $\beta\in\{1,2.5,5,10\}$ for the interaction weights, with social rewards disabled. The results in the left subplot of fig.˜4 show two key observations: (1) performance is more sensitive to the decay range $\beta$ than to the scale $\alpha$; and (2) extremely small or large $\beta$ overemphasizes very close or very distant neighbors, both of which degrade realism.
Figure 4: Influence of decay parameters $\alpha$ and $\beta$ on interaction weighting and social rewards.
Next, fixing the interaction weights at their optimal values ($\alpha=10$, $\beta=2.5$), we sweep decay scales $\alpha\in\{0.5,1,5,10\}$ and rates $\beta\in\{1,2.5,5,10\}$ for the social reward weighting. We find in the right subplot of fig.˜4 that social reward parameters have a weaker influence than interaction weights. However, very large $\alpha$ or $\beta$ make the neighborhood weights nearly uniform, which increases noise in the social reward and slightly reduces realism.
@@ 5.6 Qualitative results
We provide qualitative demonstrations of our fine-tuned model in fig.˜5, showcasing one rollout in two distinct scenarios. In both cases, the policy generates realistic multi-agent trajectories, maintaining collision-free and on-road behavior throughout the full 8-second horizon. This highlights the ability of our model to sustain expert-like motion even in complex interactive settings.
To further interpret model behavior, we visualize the per-agent rewards estimated by our decomposed discriminator, encoded as the color of each agent at every timestep. As illustrated, the discriminator consistently assigns lower rewards to agents that deviate from realistic behavior—for example, when a vehicle approaches a near-collision situation or begins drifting toward the road edge. Conversely, agents following smooth and expert-like motion patterns are assigned high rewards. Additional rollout examples, including more challenging interactive scenarios, are provided in the supplementary.
@@ Appendix A Implementation Details
We pretrain the policy network with BC for 32 epochs following CAT-K (Zhang et al., 2025a), taking about 44 hours, and then fine-tune with DecompGAIL for 2 additional epochs, taking about 10 hours. During fine-tuning, the map encoder is frozen to reduce GPU memory usage and the policy network and discriminator network share the map encoder, enabling larger batch sizes with negligible performance loss. Both pretraining and fine-tuning are conducted on 8 $\times$ H800 80GB GPUs with a total batch size of 80. We use Adam with weight decay of 0.01 for optimization, with a policy learning rate of $5\times 10^{-4}$ during pretraining and $5\times 10^{-5}$ during fine-tuning, and a discriminator learning rate of $1\times 10^{-4}$. To stabilize adversarial training, we generate fresh rollouts at each training step using the current policy, then update the discriminator and policy once. Additional hyperparameters for DecompGAIL fine-tuning are provided in table˜3.
@@
$\beta$ in $\lambda_{ij}$ 2.5
Appendix B Discriminator Gradient Penalty
To evaluate whether common gradient-penalty techniques can mitigate the instability caused by irrelevant interactions, we apply WGAN-GP (Arjovsky et al., 2017) and R1/R2 regularization (Mescheder et al., 2018) (with penalty weight fixed to 1) to both the naive PS-GAIL baseline and our DecompGAIL framework. Results are reported in table˜4.
We find that gradient penalties do modestly improve PS-GAIL, largely by constraining the discriminator’s expressiveness and thereby reducing the variance of its output scores. However, even with gradient penalties, the PS-GAIL discriminator can still be misled by irrelevant neighbor–neighbor and neighbor–map interactions, so its overall performance remains significantly below that of DecompGAIL. When adding these penalties to DecompGAIL, we observe the opposite trend: gradient penalties reduce overall realism despite slightly improving the kinematic metrics. This is expected—gradient penalties enforce smoothness in the discriminator’s mapping, but critical components of interactive and map-based realism such as collision and off-road likelihood are inherently non-smooth and depend on sharp decision boundaries. Penalizing discriminator gradients therefore harms the model’s ability to detect these events, degrading the interactive and map-based metrics.
Table 4: Effect of discriminator gradient penalties evaluated on the 2% validation split.
Model Metametric $\uparrow$ Kinematic $\uparrow$ Interactive $\uparrow$ Map-based $\uparrow$ Collision Likelihood $\uparrow$ Offroad Likelihood $\uparrow$
PS-GAIL 0.7674 0.4830 0.8012 0.8864 0.9494 0.9464
PS-GAIL + WGAN-GP 0.7723 0.5011 0.8111 0.8654 0.9538 0.9322
PS-GAIL + R1 0.7776 0.5011 0.8198 0.8814 0.9704 0.9370
PS-GAIL + R2 0.7706 0.5037 0.8091 0.8736 0.9484 0.9258
DecompGAIL + WGAN-GP 0.7865 0.5186 0.8255 0.8894 0.9796 0.9441
DecompGAIL + R1 0.7861 0.5157 0.8233 0.8928 0.9734 0.9493
DecompGAIL + R2 0.7840 0.5145 0.8212 0.8900 0.9743 0.9469
DecompGAIL 0.7889 0.5148 0.8283 0.8948 0.9837 0.9529
Appendix C Unfreezing Map Encoder
To evaluate the effect of updating the map encoder during DecompGAIL training, we conducted an additional experiment in which the map encoder is fine-tuned rather than frozen (see fig.˜6). We observe that freezing the map encoder yields slightly better training stability and marginally higher realism.
Figure 6: Effect of freezing vs. fine-tuning the map encoder. Solid lines show the mean discriminator realism scores, with shaded areas indicating standard deviation. The realism meta-metric is evaluated on the 2% validation split. Freezing the map encoder results in lower score variance and improved simulation realism throughout training.
Appendix D Ethics statement
@@
Appendix E The Use of Large Language Models (LLMs)
@@
Appendix F Reproducibility statement
We have taken multiple steps to ensure the reproducibility of our work. Our dataset and pretrained backbone are the same as those provided in the open-source CAT-K repository (https://github.com/NVlabs/catk/tree/main). Detailed descriptions of the model architecture and all hyperparameters are included in Section A and in the appendix. We also report the official performance of our method on the public WOSAC leaderboard, which enables transparent comparison against other approaches.
Appendix G Realism Score Visualization
In this section, we visualize the ego vehicle’s scene realism and interaction realism scores produced by our decomposed discriminator, for both two successful rollouts and two failure cases.
G.1 Successful Cases
We present two successful rollouts illustrating both straight-road and intersection driving and intersection. As shown in fig.˜7, the realism scores remain close to 0.5 throughout the rollout, and all agents drive without collisions or off-road behavior.
Figure 7: Straight driving case. The ego vehicle is shown with cyan edges; its color encodes the scene realism score. Surrounding vehicles are colored by their interaction realism with respect to the ego. The realism scores remain near 0.5, and all agents drive smoothly without collisions or off-road deviations.
In fig.˜8, the simulated ego vehicle chooses a turning maneuver while the ground-truth agent continues straight through the intersection, demonstrating the behavioral diversity captured by our learned traffic model. The slightly lower scene realism score in this case is likely due to low kinematic likelihood.
Figure 8: Intersection case. The ego vehicle is shown with cyan edges; its color encodes the scene realism score. Surrounding vehicles are colored by their interaction realism with respect to the ego. The simulated ego vehicle chooses a turning trajectory, whereas the ground-truth agent proceeds straight through the intersection. This illustrates the diversity of behaviors produced by our learned model. The slight decrease in scene realism mainly reflects kinematic deviations rather than unrealistic maneuvers.
G.2 Failure Cases
Figure˜9 shows a failure case in which the ego vehicle drifts outside the drivable area. As expected, our discriminator assigns a very low scene realism score at those timesteps, correctly indicating the implausible off-road behavior.
Figure 9: Off-road failure case. The ego vehicle is shown with cyan edges; its color encodes the scene realism score. Surrounding vehicles are colored by their interaction realism with respect to the ego. In the middle and right frames, the ego vehicle violates the road boundary (purple color line) and leaves the drivable surface. Our discriminator correspondingly outputs a very low scene realism score, accurately detecting the off-road behavior.
Figure˜10 presents a failure case where the ego vehicle collides with a neighboring agent. Here, our discriminator produces a very low interaction realism score, correctly flagging the implausible collision event.
Figure 10: Collision failure case. The ego vehicle is shown with cyan edges; its color encodes the scene realism score. Surrounding vehicles are colored by their interaction realism with respect to the ego. In the left frame, the ego vehicle makes contact with another vehicle. The discriminator assigns a very low interaction realism score to this agent pair, properly reflecting the implausibility of the collision.